# Exploring the mysterious high-ionization source powering [NeĀ V] in high- analog SBS0335-052Ā E with JWST/MIRI

### v1

@@ Appendix A Point source subtraction procedure
We modelled the PSF for each MIRI/MRS band using vĀ 1.4.0 WebbPSF (Perrin etĀ al. 2014) and the function calc_datacube444https://webbpsf.readthedocs.io/en/latest/jwst_ifu_datacubes.html, taking into account the same native pixel scale and spectral binning of the observed data. We created the pipeline Stage 3 products by band, selecting the option ifualign, to have them aligned with the WebbPSF models. Then, we extracted m (m for Channel 4) slices around the emission lines taken into account in this paper (see Sec.Ā 3.1) from the Stage 3 band datacubes. We localized the point source position in each slice, modeling a 2D Gaussian distribution in a region of the spectrum with solely continuum emission, where the point-source structure is more enhanced (emission lines have also extended emission). Then, using photutils PSFphotometry we found the best-fit model of the point-source emission in each slice, wavelength-by-wavelength, putting very tight constraints on its position and allowing the PSF model normalization to vary to match the flux in a region times the PSF FWHM. According to WebbPSF and JWST documentation, the PSF FWHM in Channel 1, 2, 3 and 4 is (i.e., 2, 2.5, 3, 4.5Ā px), respectively.
@@ Appendix B Other MIR diagnostic diagrams
We highlight that Fig.Ā 13 diagnostic diagram can be affected by dust attenuation since the [SĀ IV]Ā line lies in the deepest part of the silicate feature (highlighted in grey in Fig.Ā 2, Fig.Ā 3 and Fig.Ā 4). We applied a dust reddening correction (Sec.Ā 3.2) that makes the log([SĀ IV]/[NeĀ II]) line ratio Ā dex higher. Also, this diagram is unable to well-separate the low/intermediate and high-ionization emission, given that [SĀ IV]Ā and [NeĀ III]Ā have similar I.P. and behave similarly. However, it allows us to compare with models also the regions of SBSĀ 0335-052Ā E with no [NeĀ V]Ā or [OĀ IV]Ā emission, that are not shown in Fig.Ā 8 and Fig.Ā 14. The magenta line shown in Fig.Ā 13 reports the empirical relation found by Groves etĀ al. (2008), valid - with some scatter - for very different galaxies, including starbursts, ultra-luminous infrared galaxies, AGN, and also BCDs. Interestingly, SSP and SXPs models seem able to replicate the magenta line only at the highest [SĀ IV]/[NeĀ II], while shocks can only account for the lowest [SĀ IV]/[NeĀ II]Ā and [NeĀ III]/[NeĀ II]. This means that according to the explored set of models, many BCDs could be reproduced only with IMBH models, as shown for SBSĀ 0335-052Ā E.
@@ Appendix C The optical counterparts
Fig.Ā 15 bottom panels show a zoom around the HĀ and [OĀ III]Ā 4959,5007 (left) and HĀ and [NeĀ II]Ā (right) emission lines, with our best-fit shown in red and blue. To reproduce the black and spectra HĀ profiles we added up to 4 Gaussian components, with maximum widths of Ā km/s and Ā km/s, respectively. To fit the forbidden [OĀ III]Ā line we added up to 3 Gaussian components, with maximum widths of Ā km/s and Ā km/s, respectively.
The full modeling and understanding of the complex kinematics highlighted in the MUSE datacube are beyond the scope of this paper. However, the need for (even broad) multiple Gaussian components in spectra located in different parts of the galaxy and in both permitted and forbidden lines indicates that there is no BH broad line region.

### v2

@@ Appendix A Point source subtraction procedure
We modelled the PSF for each MIRI/MRS band using vĀ 1.4.0 WebbPSF (Perrin etĀ al. 2014) and the function calc_datacube777https://webbpsf.readthedocs.io/en/latest/jwst_ifu_datacubes.html, taking into account the same native pixel scale and spectral binning of the observed data. We created the pipeline Stage 3 products by band, selecting the option ifualign, to have them aligned with the WebbPSF models. Then, we extracted m (m for Channel 4) slices around the emission lines taken into account in this paper (see Sec.Ā 3.1) from the Stage 3 band datacubes. We localized the point source position in each slice, modeling a 2D Gaussian distribution in a region of the spectrum with solely continuum emission, where the point-source structure is more enhanced (emission lines have also extended emission). Then, using photutils PSFphotometry we found the best-fit model of the point-source emission in each slice, wavelength-by-wavelength, putting very tight constraints on its position and allowing the PSF model normalization to vary to match the flux in a region times the PSF FWHM. According to WebbPSF and JWST documentation, the PSF FWHM in Channel 1, 2, 3 and 4 is (i.e., 2, 2.5, 3, 4.5Ā px), respectively.
@@ Appendix B Other MIR diagnostic diagrams
We highlight that Fig.Ā 13 diagnostic diagram can be affected by dust attenuation since the [SĀ IV]Ā line lies in the deepest part of the silicate feature (highlighted in gray in Fig.Ā 2, Fig.Ā 3 and Fig.Ā 4). We applied a dust reddening correction (Sec.Ā 3.2) that makes the log([SĀ IV]/[NeĀ II]) line ratio Ā dex higher. Also, this diagram is unable to well-separate the low/intermediate and high-ionization emission, given that [SĀ IV]Ā and [NeĀ III]Ā have similar I.P. and behave similarly. However, it allows us to compare with models also the regions of SBSĀ 0335-052Ā E with no [NeĀ V]Ā or [OĀ IV]Ā emission, that are not shown in Fig.Ā 8 and Fig.Ā 14. The magenta line shown in Fig.Ā 13 reports the empirical relation found by Groves etĀ al. (2008), valid - with some scatter - for very different galaxies, including starbursts, ultra-luminous infrared galaxies, AGN, and also BCDs. Interestingly, SSP and SXPs models seem able to replicate the magenta line only at the highest [SĀ IV]/[NeĀ II], while shocks can only account for the lowest [SĀ IV]/[NeĀ II]Ā and [NeĀ III]/[NeĀ II]. This means that according to the explored set of models, many BCDs could be reproduced only with IMBH models, as shown for SBSĀ 0335-052Ā E.
@@ Appendix C The optical counterparts
Fig.Ā 15 bottom panels show a zoom around the HĀ and [OĀ III]Ā 4959,5007 (left) and HĀ and [NĀ II]Ā (right) emission lines, with our best-fit shown in red and blue. To reproduce the black and gray spectra HĀ profiles we added up to 4 Gaussian components, with maximum widths of Ā km/s and Ā km/s, respectively. To fit the forbidden [OĀ III]Ā 5007 line in the black and gray spectra we added up to 3 Gaussian components, with maximum widths of Ā km/s and Ā km/s, respectively.
The full modeling and understanding of the complex kinematics highlighted in the MUSE datacube are beyond the scope of this paper. However, the need for (even broad) multiple Gaussian components in spectra located in different parts of the galaxy and in both permitted and forbidden lines suggests that there is no BH broad line region. Similar broad components (Ā km/s) have been reported in other metal-poor galaxies (e.g., the brightest star-forming region in CGCG 007-025, del Valle-Espinosa etĀ al. 2023), whose extension also prevents its association with a BH broad line region.